OpenStudy (anonymous):

Find lone pair on central atom of this compound :- NO3-

OpenStudy (anonymous):

\[\large \bf NO_3^-\] Valence electrons in N=5+1=6(1 due to negative charge) Bond pair=3 Electron pair = (6+3)/2=9/2 Lone pair = 9/2-3

OpenStudy (anonymous):

i can't determine the lone pair on central atom ?

OpenStudy (anonymous):

@ganeshie8 @satellite73 @Preetha @pooja195 @Jesstho.-. @texaschic101 @TheSmartOne @iambatman @Kainui @freckles @dan815 @Destinymasha @arabpride @blurbendy @nincompoop

OpenStudy (anonymous):

Do you really need to tag this many people?

OpenStudy (anonymous):

yeah !... can you please help me ?

OpenStudy (anonymous):

No one will help me !!!

OpenStudy (anonymous):

@TheSmartOne can you please help me ?

syed98 (syedmohammed98):

http://prntscr.com/64wet6

syed98 (syedmohammed98):

This might help

OpenStudy (aaronq):

how do they want you to show that you've found it?

OpenStudy (aaronq):

so Nitrogen can make 5 bonds now?